OpenStudy (anonymous):

A flower was found to have two sets of opposing spirals formed by florets. The clockwise spiral has x florets. The number of florets in the counterclockwise direction is 3 more than that in the clockwise direction. These floret numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of florets in the two spirals?

OpenStudy (anonymous):

repeat?

OpenStudy (anonymous):

A. x/3=1.6 B. 3/x=1.6 C. x+3/3=1.6 D. x+3/x=1.6

OpenStudy (anonymous):

1, 1, 2, 3, 5, 8 so 5 and 8

OpenStudy (anonymous):

I just posted the answer choices?

OpenStudy (anonymous):

A. x/3=1.6